We consider the sequence of real numbers (Un) defined on N by Uo = -1, U1 = 1/2 and for every n E N, U(n+2) = U(n+1) - 1/4 Un. Where N reprents the set of natural numbers. Vn = U(n+1) - (1/2)Un.
We define the sequence (Wn) by for every n E N, (Wn) = Un / Vn.
(i) Calculate Wo and show that Wn is an arithmetic sequence and precise its common difference.
(ii) Express Wn in terms of n and calculate its limit.
1
Expert's answer
2021-01-29T00:11:18-0500

Answer:


Step 1

U0= -1

U1= "1 \over 2"


Un+2= Un+1 - "1 \over 4"Un


U2=U1- "1 \over 4"U0 = "1 \over 2"- "1 \over 4"(-1) =  "1 \over 2"+ "1 \over 4" = "3 \over 4"


U3=  U2 - "1 \over 4"U1= "3 \over 4" - "1 \over 4""1 \over 2") =  "3 \over 4"- "1 \over 8" = "5 \over 8"


U4= U3  - "1 \over 4"U2 =  "5 \over 8"- "3 \over 16" = "7 \over 16"

Now, Vn= Un+1"1 \over 2" Un


V0 = U1"1 \over 2"U0=  "1 \over 2""1 \over 2"(-1) =  "1 \over 2""1 \over 2"= 1


V1= U2"1 \over 2"U1 =  "3 \over 4"-  "1 \over 2""1 \over 2")=  "3 \over 4" - "1 \over 4" - "2 \over 4"

V2 = U3 - "1 \over 2"U2 "5 \over 8"- "1 \over 2"("3 \over 4") =  "5 \over 8" - "3 \over 8" = "2 \over 8"


V3 = U4 - "1 \over 2"U3 = "7 \over 16" -  "1 \over 2"("5 \over 8") =  "7 \over 16" - "5 \over 16"= "2 \over 16"

Step 2

i)

Now, Wn ="U_n \over V_n"

W0 = "U_0 \over V_0" = "-1 \over 1" = -1


W1= "U_1 \over V_1" = "{1 \over 2} \over { 2 \over 4}" = 1


W2= "U_2 \over V_2" = "{3 \over 4} \over { 2 \over 8}"  = 3


W3 = "U_3 \over V_3" = "{5 \over 8} \over { 2 \over 16}"  = 5


W0, W1, W2, W3, --- = -1, 1, 3, 5, ---


We can easily observe that there is a common difference of 2 in every term.


So, it forms an arithmetic sequence with d= 2 , a = -1 where d is the difference and a the initial value.


ii) Wn = 2n - 1


for limit lim  2n - 1 =

n→